# Lightweight Transformer for EEG Classification via Balanced Signed Graph Algorithm Unrolling

### v2

@@ 4.2 Self-Attention Mechanism
We review the classical self-attention mechanism in transformers (vaswani17attention, ). First, given input embedding ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ for token $i$, affinity $e_{i,j}$ between tokens $i$ and $j$ is computed as the scaled product of ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{K}},{\mathbf{Q}}\in\mathbb{R}^{E\times E}$ are the key and query matrices, respectively. Using $e_{i,j}$, non-negative and normalized attention weights $a_{i,j}$’s are computed using the softmax operator:
@@ 6.1 Experimental Setup
Datasets and settings. We evaluate our model on the Turkish Epilepsy EEG Dataset (TASCI2023252, ), which is currently the largest publicly available dataset focused on epileptic seizures. The dataset comprises 10,356 EEG recordings collected from 121 participants, including 50 patients diagnosed with generalized epilepsy and 71 healthy controls. Each recording contains 35 channels of EEG signals sampled at 500 Hz for a duration of 15 seconds. To mitigate artifacts typically observed at the beginning and end of recordings, we discard the first 2 seconds and the last 1 second.
As the default classification task setting, we follow TASCI2023252 ; LIH2023107312 ; 10558582 ; 10810442 ; Disli2025Epilepsy and divide the dataset into training, validation, and test sets in a ratio of 8: 1: 1. To assess the model’s generalization across different subjects, we also perform a leave-one-out-subject (LOSO) classification task. In this setting, data from one subject is held out as the test set, while the remaining data is used for training and validation. The training and validation sets are used for the denoising task, while the test set is reserved for classification.
@@
Baseline methods. We compare the proposed method with several competitive baselines. Non-graph-based methods include the k-Nearest Neighbors (kNN) classifier combined with Multivariate Dynamic Time Warping (MDTW) (TASCI2023252, ). The Transformer-based approach by LIH2023107312 models temporal dependencies in EEG time series data for classification. Additionally, 10558582 employs Regularized O-minus tensor network decomposition (ROD) to derive features from time-frequency (TF) maps of EEG signals. Convolutional Neural Networks (CNNs) are used by 10810442 ; Disli2025Epilepsy to extract discriminative features from EEG spectrograms, with the latter leveraging a deep CNN architecture. On the other hand, graph-based methods include DGCNN (song2018eeg, ) and GIN (zhang2021deep, ), which represent EEG data as graphs to capture complex relationships between channels, while EEGNet (Lawhern_2018, ) and FBCSPNet (8257015, ) utilize graph-based techniques to enhance feature extraction and classification performance.
@@ 6.2.1 Main Results
Method Params #
Accuracy
Precision
Recall
Specificity
F1-score
@@
MDTW + KNN (TASCI2023252, ) - 87.78 89.39 81.32 92.68 85.16
TF + ROD (10558582, ) 18,400 88.08 89.22 82.28 92.46 85.61
CWT + DCNN (Disli2025Epilepsy, ) 143,297 95.91 94.55 96.98 96.06 95.30
@@
DGCNN (song2018eeg, ) 149,466 76.74 69.56 62.74 84.60 65.97
GIN (zhang2021deep, ) 25,794 68.82 58.78 44.36 82.55 50.56
EEGNet (Lawhern_2018, ) 9,170 78.78 81.26 53.25 93.11 64.34
FBCSPNet (8257015, ) 98,242 81.76 92.80 53.40 97.67 67.79
Deep4Net (8257015, ) 321,227 78.62 73.06 64.20 86.72 68.34
@@
In the default task setting, which involves training and testing on the same dataset split, our model outperforms most baselines in terms of accuracy, precision, specificity, and F1-score, achieving 97.57%, 98.58%, 97.45%, and 98.01% respectively, with only 14,787 parameters. This demonstrates the efficiency and effectiveness of our method. While (LIH2023107312, ) based on STFT and CNN achieves a higher accuracy of 99.20% and F1-score of 99.30%, it requires several orders of magnitude more parameters (over 11.5 million), making it computationally expensive and memory-inefficient.
@@ 6.2.2 Ablation Studies
Unbalanced
Signed Graph
78.87 86.74 68.22 78.69 82.52
Balanced
Signed Graph
93.68 96.32 89.45 93.61 94.94
We also conducted further studies to assess the robustness and generalizability of our method. These include ablation studies on loss function, model architecture, and signal feature distance computation (Appendix F), validation on the TUH Abnormal EEG Corpus obeid2016temple (Appendix G), statistical significance tests for the default classification task and the LOSO task (Appendix H), and a comparison of training and inference time for graph-based baselines (Appendix I).
@@ F.1 Classification with Denoisers Trained Single MSE Loss
Loss Function
Accuracy
Precision
Recall
Specificity
F1-score
@@ F.2 Ablation Study on Temporal Sequence Length
Sequence
Length
Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@ F.3 Ablation Study on CNN Block Number
Block
Numbers
Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@ F.4 Ablation Study on Distance Metric
Distance
Metric
Accuracy
Precision
Recall
Specificity
F1-score
Convergence
Time

### v3

@@ 4.2 Self-Attention Mechanism
We review the classical self-attention mechanism in transformers (Vaswani et al., 2017). First, given input embedding ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ for token $i$, affinity $e_{i,j}$ between tokens $i$ and $j$ is computed as the scaled product of ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{K}},{\mathbf{Q}}\in\mathbb{R}^{E\times E}$ are the key and query matrices, respectively. Using $e_{i,j}$, non-negative and normalized attention weights $a_{i,j}$’s are computed using the softmax operator:
@@ 6.1 Experimental Setup
Datasets and settings. We evaluate our model on the Turkish Epilepsy EEG Dataset (Tasci et al., 2023), which is currently the largest publicly available dataset focused on epileptic seizures. The dataset comprises 10,356 EEG recordings collected from 121 participants, including 50 patients diagnosed with generalized epilepsy and 71 healthy controls. Each recording contains 35 channels of EEG signals sampled at 500 Hz for a duration of 15 seconds. To mitigate artifacts typically observed at the beginning and end of recordings, we discard the first 2 seconds and the last 1 second.
As the default classification task setting, we follow Tasci et al. (2023); Lih et al. (2023); Shen et al. (2024); Bhandage et al. (2024); Disli et al. (2025) and divide the dataset into training, validation, and test sets in a ratio of 8: 1: 1. To assess the model’s generalization across different subjects, we also perform a leave-one-out-subject (LOSO) classification task. In this setting, data from one subject is held out as the test set, while the remaining data is used for training and validation. The training and validation sets are used for the denoising task, while the test set is reserved for classification.
@@
Baseline methods. We compare the proposed method with several competitive baselines. Non-graph-based methods include the k-Nearest Neighbors (kNN) classifier combined with Multivariate Dynamic Time Warping (MDTW) (Tasci et al., 2023). The Transformer-based approach by Lih et al. (2023) models temporal dependencies in EEG time series data for classification. Additionally, Shen et al. (2024) employs Regularized O-minus tensor network decomposition (ROD) to derive features from time-frequency (TF) maps of EEG signals. Convolutional Neural Networks (CNNs) are used by Bhandage et al. (2024); Disli et al. (2025) to extract discriminative features from EEG spectrograms, with the latter leveraging a deep CNN architecture. On the other hand, graph-based methods include DGCNN (Song et al., 2018) and GIN (Zhang and Yao, 2021), which represent EEG data as graphs to capture complex relationships between channels, while EEGNet (Lawhern et al., 2018) and FBCSPNet (Schirrmeister et al., 2017) utilize graph-based techniques to enhance feature extraction and classification performance.
@@ 6.2.1 Main Results
Method Params # Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@
MDTW + KNN (Tasci et al., 2023) - 87.78 89.39 81.32 92.68 85.16
TF + ROD (Shen et al., 2024) 18,400 88.08 89.22 82.28 92.46 85.61
CWT + DCNN (Disli et al., 2025) 143,297 95.91 94.55 96.98 96.06 95.30
@@
DGCNN (Song et al., 2018) 149,466 76.74 69.56 62.74 84.60 65.97
GIN (Zhang and Yao, 2021) 25,794 68.82 58.78 44.36 82.55 50.56
EEGNet (Lawhern et al., 2018) 9,170 78.78 81.26 53.25 93.11 64.34
FBCSPNet (Schirrmeister et al., 2017) 98,242 81.76 92.80 53.40 97.67 67.79
Deep4Net (Schirrmeister et al., 2017) 321,227 78.62 73.06 64.20 86.72 68.34
@@
In the default task setting, which involves training and testing on the same dataset split, our model outperforms most baselines in terms of accuracy, precision, specificity, and F1-score, achieving 97.57%, 98.58%, 97.45%, and 98.01% respectively, with only 14,787 parameters. This demonstrates the efficiency and effectiveness of our method. While (Lih et al., 2023) based on STFT and CNN achieves a higher accuracy of 99.20% and F1-score of 99.30%, it requires several orders of magnitude more parameters (over 11.5 million), making it computationally expensive and memory-inefficient.
@@ 6.2.2 Ablation Studies
Unbalanced Signed Graph 78.87 86.74 68.22 78.69 82.52
Balanced Signed Graph 93.68 96.32 89.45 93.61 94.94
We also conducted further studies to assess the robustness and generalizability of our method. These include ablation studies on loss function, model architecture, and signal feature distance computation (Appendix F), validation on the TUH Abnormal EEG Corpus Obeid and Picone (2016) (Appendix G), statistical significance tests for the default classification task and the LOSO task (Appendix H), and a comparison of training and inference time for graph-based baselines (Appendix I).
@@ F.1 Classification with Denoisers Trained Single MSE Loss
Loss Function Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@ F.2 Ablation Study on Temporal Sequence Length
Sequence Length Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@ F.3 Ablation Study on CNN Block Number
Block Numbers Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%)
@@ F.4 Ablation Study on Distance Metric
Distance Metric Accuracy (%) Precision (%) Recall (%) Specificity (%) F1-score (%) Convergence Time